19 May 2015: Tuesday

LDAP:

  • Two issues in the LDAP update:  Two new student employees.   Found their student record and merged the employee info with it.
  • No issues in the morning ACCOUNT update.

Backups:

  • That CoM/IS system is failing again.
  • Tape clones completed.
  • Continuing to work on the check_raid for bujbod2 – got the checks to actually work and not false positive!
  • bujbod2 – nfs service does not successfully start on reboot.

Tags:

Leave a Reply

You must be logged in to post a comment.